OpenStudy (anonymous):

Solve the system of equations. 3x – 7y = 5 9y = 5x + 5

OpenStudy (precal):

do you know any methods to solve the systems?

OpenStudy (precal):

I would use the matrix and my graphing calculator if technology is allowed. I would use elimination if I had to solve it by hand.

OpenStudy (precal):

ok tell Pebbles I send my regards

OpenStudy (triciaal):

one method \[3x - 7y = 5\] \[-5x + 9y = 5\] 5 = 5 3x -7y = -5x + 9y 8x =16y x = 2y substitute in the first eqn 3(2y) -7y = 5 (6-7)y =5 y = -5 x = 2(-5) = 10 check by substituting your x and y values in either of original eqn 3(-10) -7(-5) =5 -30 + 35 = 5 5 = 5

OpenStudy (precal):

Just one note: When checking, you need to check both equations. The solutions must work for both equations.

OpenStudy (precal):

x=-10 y=-5 I doublecheck the solution with the matrix